OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8 -2y + 6 2y - 6 2y + 6 -2y - 6 @itsbribro @amistre64

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

so -x + 2y = -6 can you make x the subject of this equation...?

OpenStudy (anonymous):

yes by making y 0 and then making -x=-6

OpenStudy (campbell_st):

no, you need to retain the terms containing y.... so subtract 2y from both sides of the equation... what do you get..?

OpenStudy (anonymous):

-6-2y=-x

OpenStudy (campbell_st):

great not multiply every term by -1 to get a positive x value what do you get..?

OpenStudy (anonymous):

wait like -6(1)-2(1)=-x?

OpenStudy (campbell_st):

no.... mutliplying each term by -1 means to reverse the sign... make them all positive

OpenStudy (anonymous):

6+2y=x

OpenStudy (anonymous):

so the answer is 2y+6

OpenStudy (anonymous):

@chosenmatt